The Cheyenne Hotel in Big Sky, Montana, has accumulated records of the total electrical costs of the hotel and the number of occupancy-days over the last year. An occupancy-day represents a room rented for one day. The hotel’s business is highly seasonal, with peaks occurring during the ski season and in the summer. Month Occupancy-Days Electrical Costs January 2,830 $ 6,750 February 3,340 $ 7,410 March 3,650 $ 7,905 April 1,840 $ 5,520 May 750 $ 2,250 June 1,830 $ 5,490 July 3,890 $ 8,190 August 1,700 $ 5,100 September 4,060 $ 8,410 October 2,260 $ 6,780 November 1,350 $ 4,050 December 290 $ 870

Required: 1. Using the high-low method, estimate the fixed cost of electricity per month and the variable cost of electricity per occupancy-day. (Do not round your intermediate calculations. Round your Variable cost answer to 2 decimal places and Fixed cost element answer to nearest whole dollar amount.)

2. What other factors in addition to occupancy-days are likely to affect the variation in electrical costs from month to month? (You may select more than one answer. Single click the box with the question mark to produce a check mark for a correct answer and double click the box with the question mark to empty the box for a wrong answer. Any boxes left with a question mark will be automatically graded as incorrect.) Systematic factors like guests, switching off fans and lights. unanswered Number of days present in a month. unanswered Income taxes paid on hotel income. unchecked Seasonal factors like winter or summer. unanswered Fixed salary paid to hotel receptionist.
Answer #1
Req 1) High Low method:
Occupancy-days Electrical Costs
High 4060 8410
Low 290 870
Difference 3770 7540
Variable costs=Electrical Costs/Occupancy-days = 7540/3770 = $2 per Occupancy-day
Fixed costs = 870 - (290*2) = 290 per month
Req 2) The other factors in addition to occupancy days that will affect
the variation in electrical costs from month to month are:
a) Seasonal factors like winter or summar.